OpenStudy (anonymous):

HELPPPPPPPP (3 square root 5 + 6 square root 7)(square root 5 - square root 7)

OpenStudy (anonymous):

\[(3\sqrt{5}+6\sqrt{7})(\sqrt{5}-\sqrt{7})\]

OpenStudy (anonymous):